# Large Language Model-Based Framework for Explainable Cyberattack Detection in Automatic Generation Control Systems

### v1

@@ I Introduction
Yet, a fundamental challenge persists: the lack of transparency in ML/DL models, particularly neural networks, makes it difficult for power system operators to trust and act on their outputs. Experts may not fully trust the outputs of machine learning-based algorithms, particularly in scenarios where operational reliability is critical—a frequent requirement in the energy sector. In high-stakes environments such as AGC, where incorrect or unexplained decisions may trigger operational instability, explainability is essential for real-world adoption.
While existing explainability techniques like SHAP and LIME [5] offer insights into ML predictions, they often fall short of producing operator-friendly interpretations. Large Language Models (LLMs), in contrast, offer a unique opportunity to bridge this gap. With their ability to process domain-specific information, infer contextual meaning, and generate natural language justifications, LLMs can support human-understandable explanations of ML-based decisions—an approach increasingly explored in cybersecurity and industrial AI [6, 7].
@@
This paper introduces a hybrid framework that combines fast detection of FDIAs using lightweight ML models with explanations generated by LLMs. A key feature of the proposed framework is its alignment with power system operator workflows. Once a cyberattack is detected, the system invokes an LLM to interpret the ML model’s decision and associated signal features to produce structured, human-readable reports. These reports include estimated attack start time, affected variables, and statistical justifications—presented in operator-friendly narratives. Such reports enable operators to interpret alarms with confidence, supporting decision-making regarding mitigation, incident logging, or further investigation. By aligning with the operational needs of power system personnel and reducing the interpretability gap, the proposed framework improves decision-making, enhances operator trust and promotes actionable use of AI tools in critical power infrastructure.
@@ III-A Dataset Generation
Each sample spans 60 seconds, sampled at 0.3-second intervals, yielding 200 data points per sample—adequate for capturing system dynamics without requiring AGC oversampling. Disturbance magnitudes were drawn from a zero-mean normal distribution with a standard deviation of 0.02 p.u. Attacks were introduced post-disturbance to simulate realistic interference scenarios. Based on empirical evaluation, attack parameters $f_{f}$ and $f_{i}$ were sampled from a normal distribution centered at 0.11-0.11- 0.11 with a standard deviation of 0.02, as these values were found to yield the most effective attacks. To emulate real-world uncertainties, white Gaussian noise with zero mean and standard deviation $10^{-6}$ was added to model both process and measurement noise—representing stochastic system behavior and sensor inaccuracies, respectively. An ACE limit was enforced to constrain post-attack deviations: 0.5 for subtle attacks and 1.0 for more noticeable ones. Attacks were rescaled accordingly if they exceeded these thresholds, ensuring a wide range of difficulty levels in the generated data.
This methodology resulted in a balanced dataset of 10,000 samples—equally divided between normal and attack cases—with varied configurations. Unlike previous studies that used fixed timing, both disturbance and attack initiation times were randomized within the first 30 seconds to avoid static patterns. The simulation considered both linearized and nonlinear system behaviors, the latter incorporating generation rate constraints (GRC), deadband effects, and time delays. Figure 3 presents a sample visualization, illustrating how attack-induced deviations can be subtle, with an attack magnitude of 0.2138pu, an attack start time of 15 seconds, and $\Delta P_{\text{tie}}$ as the targeted parameter. This synthetic simulation approach was necessary due to the unavailability of real-world measurement data for security reasons. Even if such data were accessible, they would likely lack the variability needed to comprehensively evaluate FDIA detection methods—a limitation also acknowledged in prior studies that similarly relied on synthetic datasets.
@@ III-B Machine Learning Models Training
Using LLMs directly for signal classification presents challenges in both local and online deployment scenarios. LLMs demand computational resources proportional to the number of input tokens. To overcome these limitations, classification is performed by lightweight local ML models, while the LLM is employed to generate explanations based on summarized model outputs. In other words, in this work, the LLM is used as an explanation generator rather than a classifier. Following this approach, multiple ML models are evaluated, and the best-performing one is selected for the next phase. XGBoost and LightGBM are employed due to their well-established performance on structured time-series and fast inference capabilities. Random Forests (RF) is also considered for its robustness and effectiveness in similar tasks. A subset of 200 samples from each of the normal and attacked datasets is held out exclusively for LLM evaluation. From the remaining data, 70% is used for training and validation, while 30% is reserved for testing. These models are trained on signal triplets and are tasked with binary classification—labeling each sample as either normal or under FDIA.
@@ IV-B Explanation Quality Assessment
Table III compares the average performance of multiple test runs on 100 FDIA samples—selected to enable manual inspection—across three LLM variants: GPT-3.5 Turbo, GPT-4 Turbo, and GPT-4o mini. The comparison spans varying shot settings (0, 5, 10, and 20), highlighting both baseline results and the corresponding changes. The table provides insights into how each model responds to the integration of learned signal features with respect to classification accuracy, regression error, and latency. We observe inconsistency in the improvement of GPT-3.5’s accuracy as the number of shots increases. In contrast, GPT-4 Turbo exhibits relatively stable performance across all shots, with minor degradation in classification metrics and manageable variations in (Mean Absolute Error) MAE. This resilience highlights GPT-4’s robustness to additional features, likely due to its stronger architectural capacity and alignment. GPT-4o mini, however, demonstrates a more nuanced pattern with minor drops in accuracy at some shot levels Furthermore, GPT-3.5 unexpectedly exhibits smaller MAEs for both attack magnitude and time, at 10 shots, with all models exhibiting attack magnitude MAE of less than 0.02, and attack time of less than 2.2. Finally, we observe that the latency increases with the LLM’s complexity. However, the latency of all models is acceptable compared to the time scale of AGC systems, as modern AGC system collect measurement every few minutes. Recall also that attack detection is not impacted by this latency as the attack detection is performed by the ML model whose inference time is in the range of milliseconds as depicted in Table II.

### v2

@@ I Introduction
Yet, a fundamental challenge persists: the lack of transparency in ML/DL models, particularly neural networks, makes it difficult for power system operators to trust and act on their outputs. Experts may not fully trust the outputs of machine learning-based algorithms, particularly in scenarios where operational reliability is critical, a frequent requirement in the energy sector. In high-stakes environments such as AGC, where incorrect or unexplained decisions may trigger operational instability, explainability is essential for real-world adoption.
While existing explainability techniques like SHAP and LIME [5] offer insights into ML predictions, they often fall short of producing operator-friendly interpretations. Large Language Models (LLMs), in contrast, offer a unique opportunity to bridge this gap. With their ability to process domain-specific information, infer contextual meaning, and generate natural language justifications, LLMs can support human-understandable explanations of ML-based decisions, an approach increasingly explored in cybersecurity and industrial AI [6, 7].
@@
This paper introduces a hybrid framework that combines fast detection of FDIAs using lightweight ML models with explanations generated by LLMs. A key feature of the proposed framework is its alignment with power system operator workflows. Once a cyberattack is detected, the system invokes an LLM to interpret the ML model’s decision and associated signal features to produce structured, human-readable reports. These reports include estimated attack start time, affected variables, and statistical justifications, presented in operator-friendly narratives. Such reports enable operators to interpret alarms with confidence, supporting decision-making regarding mitigation, incident logging, or further investigation. By aligning with the operational needs of power system personnel and reducing the interpretability gap, the proposed framework improves decision-making, enhances operator trust and promotes actionable use of AI tools in critical power infrastructure.
@@ III-A Dataset Generation
Each sample spans 60 seconds, sampled at 0.3-second intervals, yielding 200 data points per sample, adequate for capturing system dynamics without requiring AGC oversampling. Disturbance magnitudes were drawn from a zero-mean normal distribution with a standard deviation of 0.02 p.u. Attacks were introduced post-disturbance to simulate realistic interference scenarios. Based on empirical evaluation, attack parameters $f_{f}$ and $f_{i}$ were sampled from a normal distribution centered at 0.11-0.11 with a standard deviation of 0.02, as these values were found to yield the most effective attacks. To emulate real-world uncertainties, white Gaussian noise with zero mean and standard deviation $10^{-6}$ was added to model both process and measurement noise, representing stochastic system behavior and sensor inaccuracies, respectively. An ACE limit was enforced to constrain post-attack deviations: 0.5 for subtle attacks and 1.0 for more noticeable ones. Attacks were rescaled accordingly if they exceeded these thresholds, ensuring a wide range of difficulty levels in the generated data.
This methodology resulted in a balanced dataset of 10,000 samples, equally divided between normal and attack cases, with varied configurations. Unlike previous studies that used fixed timing, both disturbance and attack initiation times were randomized within the first 30 seconds to avoid static patterns. The simulation considered both linearized and nonlinear system behaviors, the latter incorporating generation rate constraints (GRC), deadband effects, and time delays. Figure 3 presents a sample visualization, illustrating how attack-induced deviations can be subtle, with an attack magnitude of 0.2138pu, an attack start time of 15 seconds, and $\Delta P_{\text{tie}}$ as the targeted parameter. This synthetic simulation approach was necessary due to the unavailability of real-world measurement data for security reasons. Even if such data were accessible, they would likely lack the variability needed to comprehensively evaluate FDIA detection methods, a limitation also acknowledged in prior studies that similarly relied on synthetic datasets.
@@ III-B Machine Learning Models Training
Using LLMs directly for signal classification presents challenges in both local and online deployment scenarios. LLMs demand computational resources proportional to the number of input tokens. To overcome these limitations, classification is performed by lightweight local ML models, while the LLM is employed to generate explanations based on summarized model outputs. In other words, in this work, the LLM is used as an explanation generator rather than a classifier. Following this approach, multiple ML models are evaluated, and the best-performing one is selected for the next phase. XGBoost and LightGBM are employed due to their well-established performance on structured time-series and fast inference capabilities. Random Forests (RF) is also considered for its robustness and effectiveness in similar tasks. A subset of 200 samples from each of the normal and attacked datasets is held out exclusively for LLM evaluation. From the remaining data, 70% is used for training and validation, while 30% is reserved for testing. These models are trained on signal triplets and are tasked with binary classification, labeling each sample as either normal or under FDIA.
@@ III-C LLM API Setup & Prompt Structure
Figure 4: Token composition of a complete 5-shot prompt provided to the LLM
To better illustrate the structure and relative weight of each component in the constructed prompt, Fig. 4 presents the token distribution. However, the inclusion of few-shot samples introduces a noticeable overhead that increases linearly with the number of shots. Since the API is stateless, these examples must be provided with every call, as each request is processed independently. Despite this additional cost, few-shot prompting, in this case, remains more practical than fine-tuning a model, which would require specialized hardware and significantly higher computational resources for deployment.
@@ IV-B Explanation Quality Assessment
Table III compares the average performance of multiple test runs on 100 FDIA samples, selected to enable manual inspection, across three LLM variants: GPT-3.5 Turbo, GPT-4 Turbo, and GPT-4o mini. The comparison spans varying shot settings (0, 5, 10, and 20), highlighting both baseline results and the corresponding changes. The table provides insights into how each model responds to the integration of learned signal features with respect to classification accuracy, regression error, and latency. We observe inconsistency in the improvement of GPT-3.5’s accuracy as the number of shots increases. In contrast, GPT-4 Turbo exhibits relatively stable performance across all shots, with minor degradation in classification metrics and manageable variations in (Mean Absolute Error) MAE. This resilience highlights GPT-4’s robustness to additional features, likely due to its stronger architectural capacity and alignment. GPT-4o mini, however, demonstrates a more nuanced pattern with minor drops in accuracy at some shot levels Furthermore, GPT-3.5 unexpectedly exhibits smaller MAEs for both attack magnitude and time, at 10 shots, with all models exhibiting attack magnitude MAE of less than 0.02, and attack time of less than 2.2. Finally, we observe that the latency increases with the LLM’s complexity. However, the latency of all models is acceptable compared to the time scale of AGC systems, as modern AGC system collect measurement every few minutes. Recall also that attack detection is not impacted by this latency as the attack detection is performed by the ML model whose inference time is in the range of milliseconds as depicted in Table II.